OpenStudy (anonymous):

If f(x) = 2x2 - 4x + 5 and g(x) = 2x - 2, what is f(x) + g(x) ?

OpenStudy (dumbcow):

just add like terms

OpenStudy (anonymous):

dun understand

OpenStudy (dumbcow):

can you add 2x^2 with 2x ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

2x^2 -2x+ 3

OpenStudy (dumbcow):

no, they are not like terms...one has an x^2 term the other a x term the exponents on the variable must be the same for you to add them (combine)

OpenStudy (anonymous):

No, x^2& xare not similar terms!

OpenStudy (anonymous):

so answer?

OpenStudy (anonymous):

hey night will, can you help me next?

OpenStudy (anonymous):

f(x) = 2x2 - 4x + 5 and g(x) = 2x - 2 \[f(x)+g(x)=2x^2-4x+5+2x-2\] \[=x^2(2)+x(-4+2)+(5-2)\] \[=2x^2-2x+3\]

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Welcome^^